The solution in a 7 liter radiator system is 35% antifreeze. How much of the solution must be drained and replaced by a 95% antifreeze solution to obtain a 50% antifreeze solution?
Found 2 solutions by dabanfield, stanbon:
Answer by dabanfield(803)

The solution in a 7 liter radiator system is 35% antifreeze. How much of the solution must be drained and replaced by a 95% antifreeze solution to obtain a 50% antifreeze solution?
Let x be the amount to be replaced. Then we have:
.95*x + .35*(7-x) = .50*7 = 35
Solve for x.

Answer by stanbon(75887)

The solution in a 7 liter radiator system is 35% antifreeze. How much of the solution must be drained and replaced by a 95% antifreeze solution to obtain a 50% antifreeze solution?
-------------------------------------
Equation:
anti - anti + anti = anti
0.35*7 - 0.35x + 0.95x = 0.50*7
2.45 +0.60x = 3.5
0.60x = 1.05
x = 1.75 liters (amount drained and replaced)
=====================================================
Cheers,
Stan H.